What is the dimension of a rectangle with a perimeter of 460 feet?
Answer by lmeeks54(111)

There are several (actually an infinite number of) rectangles that have perimeter = 460 ft. The basis for solving this problem is understanding the equation for the perimeter of a rectangle. The perimeter of an area is the distance around the outside of that area. If your rectangle is a square (a special type of rectangle), just add up a walk around the 4 sides...
...
Let P = length of the perimeter
Let S = the length of one of the square's 4 sides
...
P = S + S + S + S = 4S
...
Consider a more 'traditional' rectangle with 2 long sides of equal length and 2 short sides of equal length:
...
Let P = perimeter
Let L = the length of the long sides
Let S = the length of the short sides
...
P = 2L + 2S
...
We're given P = 460 ft. So, any combination of 2L + 2S that totals 460 solves this problem.
...
For example, let's test this, let L = 200 and S = 30:
460 = 2(200) + 2(30)
460 = 400 + 60
460 = 460
...
Try another one, let L = 229 and S = 1:
460 = 2(229) + 2(1)
460 = 458 + 2
460 = 460
...
We could keep going on and on like this.
...
Cheers,
Lee